Find The angle between two vectors a and n as |Ã.Ñ|=Ã.Ñ. Note ~ above head denotes vector sign. || Denaoted modulus

Grade:

2 Answers

Khimraj
5 years ago
 
As per question
|A*B| = A*B 
So ||A||B|sin\Theta| = |A||B|sin\Theta
then
|sin\Theta| = sin\Theta
So \Theta \epsilon [0,\pi].
Hope it clears. If you like answer then please approve it.
Khimraj
5 years ago
As per question
|A.B| = A.B 
So ||A||B|cos\Theta| = |A||B|cos\Theta
then
|cos\Theta| = cos\Theta
So \Theta \epsilon [0,\pi/2] \cap [3\pi/2, 2\pi].
Hope it clears. If you like answer then please approve it.